OpenStudy (anonymous):

Would anyone like to explain to me how I would go about finding the domain and range of an exponential/logarithmic function?

OpenStudy (anonymous):

For example: f(x) = log(x + 1). Find the domain. I just need a little review!

OpenStudy (eyust707):

heres my trick

OpenStudy (eyust707):

"Domain, Range" x , y

OpenStudy (eyust707):

Domain is "all possible x values"

OpenStudy (eyust707):

Range is all possible y values

OpenStudy (eyust707):

Assuming it is a base 10 log... Thats the same as saying \(10^{y} = x + 1\)

OpenStudy (eyust707):

Domain would be what possible x's can you put in there and still have real numbers?

OpenStudy (eyust707):

the smallest an exponent can get is 0....

OpenStudy (eyust707):

This means that x must be greater than -1

OpenStudy (eyust707):

And we also notice that exponentials can get really big... so x in this case could get as big as you want... x would have to be less than or equal to \( \infty\)

OpenStudy (eyust707):

\[-1 \le x \le \infty\]

OpenStudy (anonymous):

@eyust707 simple way to put it! Thanks for the awesome explanation. Can you tell me anything about the purpose of asymptotes and how I would be able to find that as well?

OpenStudy (eyust707):

sure first lets look at the graph

OpenStudy (eyust707):